# EventFlow: Forecasting Temporal Point Processes with Flow Matching

### v2

@@ Abstract
Continuous-time event sequences, in which events occur at irregular intervals, are ubiquitous across a wide range of industrial and scientific domains. The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network. While autoregressive models are successful in predicting the time of a single subsequent event, their performance can degrade when forecasting longer horizons due to cascading errors and myopic predictions. We propose EventFlow, a non-autoregressive generative model for temporal point processes. The model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process. EventFlow is simple to implement and achieves a 20%-53% lower error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
@@ 1 Introduction
Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events which occur at random times. Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley & Vere-Jones, 2003) which characterizes the distribution over such sequences. Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
@@
While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham & Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences. This has motivated the use of neural networks (Du et al., 2016; Mei & Eisner, 2017) in modeling TPPs. To date, most neural network based TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020), where a model is trained to predict the next event time given an observed history of events. However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow. Despite their popularity, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
@@
Moreover, existing models are typically trained via a maximum likelihood procedure (see Section 3) which involves computing the CDF implied by the learned model. When using a neural model, computing this CDF necessitates techniques such as Monte Carlo estimation to properly compute the loss (Mei & Eisner, 2017). In addition, sampling from intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Yang et al., 2022) is nontrivial, requiring an expensive and difficult to implement approach based on the thinning algorithm (Lewis & Shedler, 1979; Ogata, 1981; Xue et al., 2024).
Motivated by these limitations, we propose EventFlow, a generative model which directly learns the joint event time distributions, thus allowing us to avoid autoregressive sampling altogether. Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs, where we learn a continuous flow from a reference TPP to our data TPP. At an intuitive level, samples from our model are generated by drawing a collection of event times from a reference distribution and flowing these events along a learned vector field. The number of events is fixed throughout this process, decoupling the event counts and their times, so that the distribution over event counts can be learned or otherwise specified. See Figure 1 for an illustration. More specifically,
  • We propose EventFlow, a novel generative model for temporal point processes. Our model is suitable for both unconditional generation tasks (i.e., generating draws from the underlying data TPP) and conditional generation tasks (e.g., forecasting future events given a history), and is able to forecast multiple events simultaneously.
  • We provide new perspective on modeling TPPs which sidesteps common pitfalls. In particular, the key idea of EventFlow is to decompose the generative process into a learned event count distribution and a flow-based generative model for the joint distribution of event times. Our model is non-autoregressive, easy to sample from, and straightforward to implement.
  • On standard benchmarks, we demonstrate that EventFlow obtains state-of-the-art performance on a multi-step forecasting task and matches or exceeds the performance of state-of-the-art models for unconditional generation. In particular, we achieve a 20%-53% reduction in forecasting error from the nearest baseline. Compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass.
@@
Figure 1: All illustration of forecasting with our EventFlow method. The horizontal axis indicates the flow time $s$, and the vertical axis indicates the support of the TPP $\mathcal{T}=[0,T]$. We first encode the observed history $\mathcal{H}$ into an embedding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$. At $s=0$, we independently draw $n$ events in the forecasting window $[T_{0},T_{0}+\Delta T]$ from a fixed reference distribution, constituting a sample $\gamma_{0}$ from a mixed-binomial TPP. Each event can be thought of as a particle, which is assigned a velocity by a neural network $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$. Each particle flows along its corresponding velocity field until reaching its terminal point at $s=1$, whereby we obtain a forecasted sequence $\gamma_{1}$.
@@ 2 Related Work
The statistical modeling of temporal point processes (TPPs) is a classical subject with a long history (Daley & Vere-Jones, 2003; Hawkes, 1971; Isham & Westcott, 1979). The contemporary modeling paradigm, based on neural networks (Du et al., 2016), typically operates by learning a history encoder and an event decoder. The history encoder seeks to learn a fixed-dimensional vector representation of the history of a sequence up to some given time, and the decoder seeks to model a distribution over the subsequent event(s).
Numerous models have been proposed for both components. Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022). While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead.
Similarly, a wide range of forms for the event decoder have also been proposed. The most common approach is to parametrize a conditional intensity function via a neural network (Mei & Eisner, 2017; Zuo et al., 2020; Du et al., 2016). An alternative approach is to use generative models as decoders. These models often do not assume a parametric form for the decoder, enhancing their flexibility. For instance, Xiao et al. (2017b) propose the use of W-GANs to generate new events. Similarly, Shchur et al. (2020a) learn the distribution over the next inter-arrival time via a normalizing flow. Lin et al. (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs. Despite the flexibility of these models, these approaches are all autoregressive in nature, making them ill-suited for multi-step forecasting tasks. In contrast, Lüdke et al. (2023) propose a diffusion-style model which is able to avoid autoregressive sampling via an iterative refinement procedure.
Our work can be viewed as a novel approach to building flexible decoders for TPPs, extending flow matching to the setting of continuous-time event sequences. In contrast previous generative approaches, our model is trained in a likelihood-free manner and is non-autoregressive. The work of Lüdke et al. (2023) is most closely related to ours, but we emphasize that Lüdke et al. (2023) require an involved training and sampling procedure. In contrast, our method is straightforward to both implement and sample from, while simultaneously outperforming existing approaches.
@@
The recently introduced flow matching framework (or stochastic interpolants) (Lipman et al., 2023; Albergo & Vanden-Eijnden, 2023; Liu et al., 2023) describes a class of generative models which are closely related to both normalizing flows (Papamakarios et al., 2021) and diffusion models (Ho et al., 2020; Song et al., 2021). Intuitively, these models learn a path of probability distributions which interpolates between a fixed reference distribution and the data distribution. These models are a popular alternative to diffusion, providing greater flexibility in model design, with recent applications in image generation (Ma et al., 2024; Dao et al., 2023), DNA and protein design (Stark et al., 2024; Campbell et al., 2024), and point cloud generation (Buhmann et al., 2023; Wu et al., 2023). To the best of our knowledge, our work is the first to explore flow matching for TPPs.
@@ 3 Autoregressive TPP Models
We first provide a brief review of autoregressive point process models and discuss their shortcomings. Informally, one may think of an event sequence as a set $\{t_{k}\}_{k=1}^{n}$ of increasing event times. We will use $\mathcal{H}_{t}$ to represent the history of a sample up to (and including) time $t$, i.e., $\mathcal{H}_{t}=\{t_{k}:t_{k}\leq t\}$. Similarly, we use $\mathcal{H}_{t^{-}}=\{t_{k}:t_{k}<t\}$ to represent the history of a sample prior to time $t$. In the autoregressive setting, the time of a single future event $t$ is modeled conditioned on the observed history of a sequence. This is typically achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
In the first approach, a conditional probability density of the form $p_{n}(t\mid\mathcal{H}_{t_{n}})$ is learned, allowing us to specify a joint distribution over event times $p(t_{1},\dots,t_{n})$ autoregressively via ${p_{n}(t_{1},\dots,t_{n})=p(t_{1})\prod_{k=2}^{n}p(t_{k}\mid\mathcal{H}_{t_{k-1}})}$. Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t_{n}})/(1-F(t\mid\mathcal{H}_{t_{n}}))$, where $F(t\mid\mathcal{H}_{t_{n}})=\int_{t_{n}}^{t}p(s\mid\mathcal{H}_{t_{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t_{n}})$. Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the previous $n$ events and given that no events have occurred since $t_{n}$. By integrating $\lambda^{*}(t)$, one can show that
$F(t\mid\mathcal{H}_{t_{n}})=1-\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)\qquad p(t\mid\mathcal{H}_{t_{n}})=\lambda^{*}(t)\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (1)
@@
Suppose we observe an event sequence $\{t_{k}\}_{k=1}^{n}$ on the interval $[0,T]$. The likelihood of this sequence can be seen loosely as the probability of seeing precisely $n$ events at these times. The likelihood may be expressed in terms of either the density or intensity via
$L\left(\{t_{k}\}_{k=1}^{n}\right)=p_{n}(t_{1},\dots,t_{n})\left(1-F(T\mid\mathcal{H}_{t_{n}})\right)=\left(\prod_{k=1}^{n}\lambda^{*}(t_{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (2)
where the CDF term is included to indicate that no events beyond $t_{n}$ have occurred and $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity. Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei & Eisner, 2017; Shchur et al., 2020a). We emphasize that this likelihood is not simply the joint event-time density $p_{n}(t_{1},\dots,t_{n})$, as the likelihood also measures the fact that no events occur between $t_{n}$ and $T$.
It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice. For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei & Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$. Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (2). Moreover, some approaches, such as the diffusion-based approach of Lin et al. (2022), are only trained to maximize an ELBO of $p(t_{k}\mid t_{1},\dots,t_{k-1})$, and are thus unable to compute the proper likelihood in Equation (2).
@@
In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences. For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions. However, sampling from existing models can be difficult. The flow-based model of Shchur et al. (2020a) requires a numerical approximation to the inverse of the model to perform sampling. The diffusion-based approach of Lin et al. (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences. Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023). This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei & Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (1) is prohibitively expensive. Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis & Shedler, 1979). However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024). For instance, the thinning algorithm requires one to know an upper bound on the intensity, which is in practice selected heuristically.

### v3

@@ Abstract
Continuous-time event sequences, in which events occur at irregular intervals, are ubiquitous across a wide range of industrial and scientific domains. The contemporary modeling paradigm is to treat such data as realizations of a temporal point process, and in machine learning it is common to model temporal point processes in an autoregressive fashion using a neural network. While autoregressive models are successful in predicting the time of a single subsequent event, their performance can degrade when forecasting longer horizons due to cascading errors and myopic predictions. We propose EventFlow, a non-autoregressive generative model for temporal point processes. The model builds on the flow matching framework in order to directly learn joint distributions over event times, side-stepping the autoregressive process. EventFlow is simple to implement and achieves a 20%-53% lower forecast error than the nearest baseline on standard TPP benchmarks while simultaneously using fewer model calls at sampling time.
@@ 1 Introduction
Many stochastic processes, ranging from the occurrence of earthquakes (Ogata, 1998) to consumer behavior (Xu et al., 2014), are best understood as a sequence of discrete events that occur at random times. Any observed event sequence, consisting of one or more event times, may be viewed as a draw from a temporal point process (TPP) (Daley and Vere-Jones, 2003) which characterizes the distribution over such sequences. Given a collection of observed event sequences, faithfully modeling the underlying TPP is critical in both understanding and forecasting the phenomenon of interest.
@@
While multiple different parametric TPP models have been proposed (Hawkes, 1971; Isham and Westcott, 1979), their limited flexibility limits their application when modeling complex real-world sequences. This has motivated the use of neural networks (Du et al., 2016; Mei and Eisner, 2017) in modeling TPPs. To date, most neural TPP models are autoregressive in nature (Shchur et al., 2020a; Zhang et al., 2020; Shchur et al., 2021), where a model is trained to predict only a single subsequent event time given an observed history of events. However, in many tasks, we are interested not only in the next event, but in the entire sequence of events which is to follow. While autoregressive neural TPP models can be applied in this setting, their performance in many-step forecasting tasks can be unsatisfactory due to compounding errors arising from the autoregressive sampling procedure (Xue et al., 2022; Lüdke et al., 2023).
@@
Moreover, existing models are typically trained via a maximum likelihood procedure (see Section 3) which involves computing the CDF implied by the learned model. When using a neural model, computing this CDF often requires techniques such as Monte Carlo estimation to compute the loss (Mei and Eisner, 2017). In addition, sampling from intensity-based models (Du et al., 2016; Mei and Eisner, 2017; Yang et al., 2022) is nontrivial, requiring an expensive and difficult to implement approach based on the thinning algorithm (Lewis and Shedler, 1979; Ogata, 1981; Xue et al., 2024).
@@
Figure 1: All illustration of forecasting with our EventFlow method. The horizontal axis indicates the flow time $s$, and the vertical axis indicates the support of the TPP $\mathcal{T}=[0,T]$. We first encode the observed history $\mathcal{H}$ into an embedding $e_{\mathcal{H}}=f_{\theta}(\mathcal{H})$. At $s=0$, we independently draw $n$ events in the forecasting window $T_{0},T_{0}+\Delta T]$ from a fixed reference distribution, constituting a sample $\gamma_{0}$ from a mixed-binomial TPP. Each event can be thought of as a particle, which is assigned a velocity by a neural network $v_{\theta}(\gamma_{s},s,e_{\mathcal{H}})$. Each particle flows along its velocity field until reaching its terminal point at $s=1$, whereby we obtain a forecasted sequence $\gamma_{1}$.
Motivated by these limitations, we propose EventFlow, a generative model which directly learns the full joint distribution over future event times, allowing us to avoid autoregressive sampling altogether. Our proposed model extends the flow matching framework (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) to the setting of TPPs. To generate a sample, we first draw a set of random event times from a simple reference distribution and then transport them through a learned vector field to obtain a sequence of predicted times. The number of events is held fixed during this transformation, decoupling the modeling of event counts from that of event times. See Figure 1 for an illustration.
This yields a flexible formulation that can be used for both unconditional generation (sampling from the underlying TPP) and conditional generation (forecasting future events given a history). More broadly, EventFlow provides an alternative perspective on TPP modeling: rather than specifying an intensity function and relying on sequential simulation, we decompose the generative process into a distribution over event counts and a flow-based model over event times. This results in a non-autoregressive model that is straightforward to train and efficient to sample from.
Empirically, we demonstrate that EventFlow obtains state-of-the-art performance on multi-step forecasting benchmarks and is competitive with leading approaches for unconditional generation. In particular, our model reduces forecast error by 20%–53% relative to the strongest baselines. Further, compared with existing methods, which can require hundreds of model evaluations at sampling time, we demonstrate that EventFlow can achieve competitive performance with only a single forward pass of the underlying neural network.
@@ 2 Related Work
The statistical modeling of temporal point processes (TPPs) has a long history (Daley and Vere-Jones, 2003; Hawkes, 1971; Isham and Westcott, 1979). The contemporary modeling paradigm, based on neural networks (Du et al., 2016; Shchur et al., 2021), typically operates by learning a history encoder and an event decoder. The history encoder seeks to learn a fixed-dimensional vector representation of the sequence of events which have been observed or already predicted, and the decoder seeks to model a distribution over the subsequent event time(s).
Popular choices for the history encoder include RNN-based models (Du et al., 2016; Shchur et al., 2020a; Mei et al., 2019) or attention-based models (Zhang et al., 2020; Zuo et al., 2020; Yang et al., 2022). While attention-based encoders can provide longer-range contexts, this benefit typically comes at the cost of additional memory overhead.
For the event decoder, a common approach is to parametrize a conditional intensity function using a neural network that takes as input a learned representation of the event history (Du et al., 2016; Mei and Eisner, 2017; Zuo et al., 2020). An alternative and more recent approach is to use generative models as decoders. These models often do not assume a parametric form for the decoder, enhancing their flexibility. For instance, Xiao et al. (2017b) propose the use of W-GANs to generate new events. Similarly, Shchur et al. (2020a) learn a distribution over the next event time via a normalizing flow. Lin et al. (2022) benchmark several choices of generative models, including diffusion, GANs, and VAEs. Regardless of the specific choice of decoder, these approaches are all autoregressive, making them ill-suited for multi-step forecasting tasks. Notable exceptions are the works of Lüdke et al. (2023, 2026), which propose flow and diffusion based models that avoids autoregressive sampling through an iterative refinement procedure.
Our EventFlow model can be viewed as a flexible, non-autoregressive decoder for TPPs, obtained by extending flow matching to continuous-time event sequences. Among existing approaches, Lüdke et al. (2023, 2026) are most closely related in spirit. However, their method relies on a relatively involved training and sampling pipeline based on iterative refinement. In contrast, EventFlow admits a particularly simple regression-based formulation which is straightforward to implement. While previous works have studied marked TPPs (Du et al., 2016; Mei and Eisner, 2017; Draxler et al., 2025; Chang et al., 2025), EventFlow is focused on modeling the event times themselves.
@@
The flow matching framework (or stochastic interpolants) (Lipman et al., 2023, 2024; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) describes a class of generative models which are closely related to normalizing flows (Papamakarios et al., 2021). Flow matching has been explored for applications including image generation (Dao et al., 2023; Ma et al., 2024), DNA and protein design (Stark et al., 2024; Campbell et al., 2024), and 3D modeling (Buhmann et al., 2023; Wu et al., 2023; Xiang et al., 2025), but our work is the first to apply flow matching for TPPs.
@@ 3 Autoregressive TPP Models
We first provide a brief, informal review of autoregressive point process models and discuss their shortcomings. Informally, an event sequence is a set $\{t^{k}\}_{k=1}^{n}$ of increasing event times. We will use $\mathcal{H}_{t}$ to represent the history of a sample up to (and including) time $t$, i.e., $\mathcal{H}_{t}=\{t^{k}:t^{k}\leq t\}$. Similarly, $\mathcal{H}_{t^{-}}=\{t^{k}:t^{k}<t\}$ represents the history prior to time $t$. In the autoregressive setting, the time of a single future event $t$ is modeled conditioned on the observed history of a sequence. This is often achieved by either directly modeling a distribution over $t$ (Shchur et al., 2020a), or equivalently by modeling a conditional intensity function (Du et al., 2016).
More precisely, one models conditional densities of the form $p(t^{k}\mid\mathcal{H}_{t^{k-1}})$, which describe the distribution of the next event time given the history up to the previous event. This induces a joint distribution over event times $p(t^{1},\dots,t^{n})$ autoregressively via $p(t^{1},\dots,t^{n})=p(t^{1})\prod_{k=2}^{n}p(t^{k}\mid\mathcal{H}_{t^{k-1}})$}. Alternatively, we may define the conditional intensity $\lambda^{*}(t):=\lambda(t\mid\mathcal{H}_{t^{-}})=p(t\mid\mathcal{H}_{t^{n}})/(1-F(t\mid\mathcal{H}_{t^{n}}))$, where $F(t\mid\mathcal{H}_{t^{n}})=\int_{t^{n}}^{t}p(s\mid\mathcal{H}_{t^{n}})\,\mathrm{d}s$ is the CDF associated with $p(t\mid\mathcal{H}_{t^{n}})$. Informally, the conditional intensity $\lambda^{*}(t)$ can be thought of (Rasmussen, 2011) as the instantaneous rate of occurrence of events at time $t$ given the history up to $t^{n}$ and that no events have occurred in $(t^{n},t)$. By integrating $\lambda^{*}(t)$, one can verify
$F(t\mid\mathcal{H}_{t^{n}})=1-\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (1)
$p(t\mid\mathcal{H}_{t^{n}})=\lambda^{*}(t)\exp\left(-\int_{t_{n}}^{t}\lambda^{*}(s)\,\mathrm{d}s\right)$ (2)
@@
Suppose we observe an event sequence $\{t^{k}\}_{k=1}^{n}$ on the interval $[0,T]$. The likelihood of this sequence can be seen loosely as the probability of observing events at these times and no others within the observation window. The likelihood may be expressed in terms of either the density or intensity via
$L\left(\{t^{k}\}_{k=1}^{n}\right)$ $=p(t^{1},\dots,t^{n})\left(1-F(T\mid\mathcal{H}_{t^{n}})\right)$ (3)
$=\left(\prod_{k=1}^{n}\lambda^{*}(t^{k})\right)\exp\left(-\Lambda^{*}(T)\right)$ (4)
where $\Lambda^{*}(T)=\int_{0}^{T}\lambda^{*}(s)\,\mathrm{d}s$ is the total intensity. The survival term $1-F(T\mid\mathcal{H}_{t^{n}})$ accounts for the absence of events in $(t^{n},T]$. Autoregressive models are typically trained by maximizing this likelihood (Du et al., 2016; Mei and Eisner, 2017; Shchur et al., 2020a). Critically, we emphasize that $p(t^{1},\dots,t^{n})$ captures a joint density over $n$ event times, but does not account for what occurs after the final event $t^{n}$. In contrast, the likelihood $L\left(\{t^{k}\}_{k=1}^{n}\right)$ corresponds to observations in a finite window $[0,T]$, and therefore includes an additional CDF term to account for the fact that no events were observed in $(t^{n},T]$.
It is worth noting that evaluating $L(\{t_{k}\})$ can be non-trivial in practice. For models that parametrize $\lambda^{*}(t)$ via a neural network (Du et al., 2016; Mei and Eisner, 2017), computing the total intensity $\Lambda^{*}(T)$ is often done via a Monte Carlo integral, requiring many forward passes of the model to evaluate $\lambda^{*}(t)$ at different values of $t$. Models which directly parametrize the density $p(t\mid\mathcal{H}_{t})$ suffer from the same drawback when computing the corresponding CDF in Equation (3). Moreover, some approaches, such as the diffusion-based approach of Lin et al. (2022), are only trained to maximize an ELBO of $p(t_{k}\mid t_{1},\dots,t_{k-1})$, and are thus unable to compute the proper likelihood in Equation (3).
@@
In many tasks, we are interested not only in an accurate model of the intensity (or distribution), but also sampling new event sequences. For instance, when forecasting an event sequence, we may want to generate several forecasts in order to provide uncertainty quantification over these predictions. However, sampling from existing models can be difficult. The flow-based model of Shchur et al. (2020a) requires a numerical approximation to the inverse of the model to perform sampling. The diffusion-based approach of Lin et al. (2022) can require several hundred forward passes of the model to generate a single event time, rendering it costly when generating long sequences. Moreover, the predictive performance of autoregressive models is often unsatisfactory on multi-step generation tasks due to the accumulation of errors over many steps (Lin et al., 2021; Lüdke et al., 2023). This difficulty is particularly pronounced for intensity-based models (Du et al., 2016; Mei and Eisner, 2017; Zhang et al., 2020), where naively computing the implied distribution in Equation (2) is prohibitively expensive. Instead, sampling from intensity-based models is typically achieved via the thinning algorithm (Ogata, 1981; Lewis and Shedler, 1979). However, this algorithm has several hyperparameters to tune, is challenging to parallelize, and can be difficult for practitioners to implement (Xue et al., 2024). For instance, the thinning algorithm requires one to know an upper bound on the intensity, which is in practice selected heuristically.